OpenStudy (anonymous):

find f'(-3) if f(x)=x^3/5-10x

OpenStudy (anonymous):

-27/5+30 -27/35

OpenStudy (anonymous):

-.7714

OpenStudy (anonymous):

Differentiate to get: \[f'(x)=\frac{3x^2}{5}-10\] Then substitute -3 in for x: \[f'(-3)=\frac{3(-3)^2}{5}-10=\frac{27}{5}-10=-4 \frac{3}{5}\]